Refer to the following information for Questions 1, 2 and 3.
A box contains 10 chips. The chips are numbered 1 through 10. Otherwise, the chips are identical. From this box, we draw one chip at random, and record its value. We then put the chip back in the box. We repeat this process two more times, making three draws in all from this box.
1. How many elements are in the sample space of this experiment?
2. What is the probability that the three numbers drawn are not all the same?
2. What is the probability that the three numbers drawn are all different?

Answer by stanbon(75887)

A box contains 10 chips. The chips are numbered 1 through 10. Otherwise, the chips are identical. From this box, we draw one chip at random, and record its value. We then put the chip back in the box. We repeat this process two more times, making three draws in all from this box.
1. How many elements are in the sample space of this experiment?
Ans: 10^3
------------------------
2. What is the probability that the three numbers drawn are not all the same?
P(all three ARE the same)= 10(1/10)^3 = 1/100
P(all are NOT eht same) = 99/100
---------------------------------------------
2. What is the probability that the three numbers drawn are all different?
Ans:
1st draw: any number
2nd draw: 1 of 9 numbers
3rd draw: 1 of 8 numbers
P(all are different): 1*(1/9)(1/8) = 1/72
----
Cheers,
Stan H.